At 1 bar, how much energy is required to heat 37.0 g of H2O(s) at –12.0 °C to H,O(g) at 119.0 °C? 9 =
Answer #1

Ti = -12.0 oC
Tf = 119.0 oC
here
Cs = 2.09 J/goC

Heat required to convert solid from -12.0 oC to 0.0 oC
Q1 = m*Cs*(Tf-Ti)
= 37 g * 2.09 J/goC *(0--12) oC
= 927.96 J

Lf = 333.6 J/g

Heat required to convert solid to liquid at 0.0 oC
Q2 = m*Lf
= 37.0g *333.6 J/g
= 12343.2 J

Cl = 4.184 J/goC

Heat required to convert liquid from 0.0 oC to 100.0 oC
Q3 = m*Cl*(Tf-Ti)
= 37 g * 4.184 J/goC *(100-0) oC
= 15480.8 J

Lv = 2260.0 J/g

Heat required to convert liquid to gas at 100.0 oC
Q4 = m*Lv
= 37.0g *2260.0 J/g
= 83620 J

Cg = 2.01 J/goC

Heat required to convert vapour from 100.0 oC to 119.0 oC
Q5 = m*Cg*(Tf-Ti)
= 37 g * 2.01 J/goC *(119-100) oC
= 1413.03 J

Total heat required = Q1 + Q2 + Q3 + Q4 + Q5
= 927.96 J + 12343.2 J + 15480.8 J + 83620 J + 1413.03 J
= 113785 J
= 113.8 KJ
Answer: 113.8 KJ